Define and calculate the normal and abnormal Zeeman effect for the n = 1 and n = 2 states of the hydrogen atom.
Answer #1

DATE: 1 Normal! The normal zeeman effect which is - shown by all lines due to transmission (between singlet (s=o) states of aof electron [2] Abnormal / anomalous- Its dependent on the spin MJ = -J -- tJ Ja ī+ Ź for nat L=0, s= 1/2 So, My = + / - 1 SuDATE: 기2 / ~/2 - 3/3 Perk + (2 /EA. YT- F(A, m₂ = -1/2 + + 2 OXY T- IT I | AM = 0,